OpenStudy (anonymous):

In isosceles Symbol for triangle A B C., B is the vertex. The measure of Symbol for angle B. can be represented as (2x+12). The measure of Symbol for angle A. can be represented as (4x-1). Find the measure of all three angles of the triangle. Be sure to name the angle with it’s measurement (ex: angle A= ). In isosceles Symbol for triangle A B C., B is the vertex. The measure of Symbol for angle B. can be represented as (2x+12). The measure of Symbol for angle A. can be represented as (4x-1). Find the measure of all three angles of the triangle. Be sure to name the angle with it’s measurement (ex: angle A= ). @Mathematics

OpenStudy (turingtest):

is is isosceles so the two base angles are the same (C=A). This means that 2x+12+2(4x-1)=180 10x+10=180 10x=170 x=17 A=C=4(17)-1=67 B=2(17)+12=46

OpenStudy (turingtest):

It^^^

OpenStudy (anonymous):

:DD thank uu

OpenStudy (anonymous):

A = 67, B = 46 and C=67

OpenStudy (anonymous):

<3